# High-Dimensional Sparse Clustering via Iterative Semidefinite Programming Relaxed K-Means

### v1

@@ 2.2 Minimax rate for separation bound in isotropic case
and $\mathcal{H}(K):=\bigl{\{}\mathbf{H}=(H_{i,k})\in\{0,1\}^{n\times K}:\sum_{k=1}^{n}H_{i,k}=1~{}\text{for all }i=1,\ldots,n\bigr{\}}$. For a given clustering algorithm $\hat{\mathbf{H}}:\mathbb{R}^{n}\to\mathcal{H}(K)$, we assess its performance using the worst-case probability of exact cluster recovery over the distribution class $\Theta(s,\Delta,K)$:
$\sup_{\boldsymbol{\theta}\in\Theta(s,\Delta,K)}\min_{\pi\in S_{K}}\mathbb{P}_{\boldsymbol{\theta}}\bigl{(}\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H}\bigr{)},$ (5)
where $S_{K}$ is the set of label permutations of $[K]$. For brevity, we omit $\min_{\pi\in S_{K}}$ henceforth.
Given a problem instance $\boldsymbol{\theta}^{\ast}=(\boldsymbol{\mu}_{1}^{\ast},\ldots,\boldsymbol{\mu}_{k}^{\ast},\mathbf{H}^{\ast})\in\Theta(s,\Delta,K)$, the true cluster assignment $\mathbf{H}^{\ast}$ has one-to-one correspondence to the block-diagonal matrix $\mathbf{Z}^{\ast}:=\mathbf{H}^{\ast}\mathbf{B}^{\ast}(\mathbf{H}^{\ast})^{\top}=\sum_{k=1}^{K}|G_{k}^{\ast}|^{-1}\mathbf{1}_{G_{k}^{\ast}}\mathbf{1}_{G_{k}^{\ast}}^{\top},$ up to label permutations. Let us define:
$m$ $:=2\min_{1\leq k\neq l\leq K}\left\{\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}\right\},\quad\Delta_{S\cap S_{0}}^{2}:=\min_{1\leq k\neq l\leq K}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2},~{}\text{and}$
$\mathcal{S}$ :=$:=\left\{S:S\subset[p],|S|\leq\sqrt{p},\Delta_{S\cap S_{0}}^{2}\gtrsim\sigma^{2}\left(\log n+\frac{|S|\log p}{m}+\sqrt{\frac{|S|\log p}{m}}\right)\right\}.$ (6)
The following theorem presents a separation condition for exact recovery by the SDP K-means.
Theorem 1 (Separation condition).
Assume that there exists a universal constant $C_{1}$ such that $m\geq C_{1}n/\log n.$ Also assume $|G_{k}^{\ast}|\geq 2$ for all $k\in[K]$. Let $\hat{\mathbf{Z}}(S)$ be the minimizer of (3) corresponding to the subset $S\in\mathcal{S}$. Then $\mathbb{P}(\hat{\mathbf{Z}}(S)=\mathbf{Z}^{*},\forall S\in\mathcal{S})\geq 1-C_{3}K/n,$ where $C_{3}>0$ is some constant.
The proof of Theorem 1 is presented in Appendix C. The stated probability in Theorem 1 holds uniformly over all support sets in $\mathcal{S}$, ensuring the validity of iterative algorithms that solve multiple SDP problems over varying subsets $S$ across iterations. The assumption on $m$ prevents overly imbalanced clusters. It is essential for theoretical guarantees for clustering and is either explicitly stated (Azizyan et al.,, 2013; Cai et al.,, 2019; Chen and Yang,, 2021; Chen and Zhang,, 2024; Jin and Wang,, 2016; Löffler et al.,, 2021; Verzelen and Arias-Castro,, 2017) or implicitly acknowledged (Giraud and Verzelen,, 2019; Royer,, 2017) in much of the literature.
Next, we investigate the fundamental limits of sparse Gaussian mixture clustering. Specifically, we establish that for any clustering method to guarantee worst-case exact recovery over the $\ell_{2}$-separated class in (4), its required separation threshold must be at least $\sigma^{2}\log n.$
@@
Consider the regime $(s\log p)/n=o(1)$. For $n\geq 2$, no clustering rule can guarantee exact recovery in the worst case over the class $\overline{\Theta}:=\Theta\bigl{(}s,C_{4}\sigma^{2}\log n,K\bigr{)},$ with $C_{4}$ some small positive constant. In other words, we have $\inf_{\hat{\mathbf{H}}}\sup_{\boldsymbol{\theta}\in\overline{\Theta}}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H})\geq C_{5},$ with $C_{5}$ a constant.
The proof of Theorem2 is presented in Appendix D. Theorem 2 implies that, under the regime of $|S|\log p/m\lesssim\log n$, the separation condition in Theorem 1 (i.e., lower bound for $\Delta^{2}_{S\cap S_{0}}$) is dominated by $\log n$ and thus matches the bound in Theorem 2 up to a constant.
The above theoretical results also indicate that, in the sparse regime where $|S_{0}|\ll p$, variable selection can be critical for clustering accuracy. To understand that, consider the simple scenario where $K=2$, $|G_{1}^{\ast}|=|G_{2}^{\ast}|$ and $(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})_{S_{0}}=\mu_{0}\mathbf{1}_{|S_{0}|}$. Then Theorems 1 and 2 prove that for any set $S$ satisfying $|S|\lesssim\sqrt{p}$, and $|S|\lesssim(n\log n)/\log p$ , exact recovery is possible if and only if
@@
This condition becomes very stringent if $S$ is a very poor estimate of $S_{0}$ (i.e., missing too many signal variables). It also suggests that a slight overestimate of $S_{0}$ may not be very damaging. These motivate us to design an algorithm that incorporates some conservative variable selection method into clustering.
3 Proposed iterative algorithms
This section presents our proposed algorithms under two scenarios. Section 3.1 considers the case where the covariance matrix is known, which includes the setting in Section 2.2 as a special case with $\boldsymbol{\Sigma}^{\ast}=\sigma^{2}\mathbf{I}_{p}$. Section 3.2 then extends the methodology to scenarios with unknown covariance. For simplicity, the algorithms are presented for $K=2$, though they naturally extend to any $K$.
3.1 Iterative algorithm under known covariance using the innovated transformation
The success of feature selection critically depends on the signal strength. In the related but different problem of classification, it has been explored in the literature Fan et al., (2013, 2015) that the innovated data transformation of $\tilde{\mathbf{X}}:=\boldsymbol{\Omega}^{\ast}\mathbf{X}$ boosts the signal strength for important features the most for a general precision matrix that may not be identity. See Fan et al., (2013) for detailed arguments and optimality study for the innovated transformation in Gaussian classification.
Our proposed procedure works with the transformed data $\tilde{\mathbf{X}}$ (or its estimate when $\boldsymbol{\Omega}^{\ast}$ unknown, in Section 3.2). It alternates between a variable selection step and a clustering step, detailed below, with the complete procedure summarized in Algorithm 2.
Variable selection step.
As argued in Cai et al., (2019), clustering is closely related to classification where the optimal rule is the Bayes rule with linear classification boundary $\boldsymbol{x}^{T}\boldsymbol{\beta}^{\ast}$ for any given test point $\boldsymbol{x}\in\mathbb{R}^{p}$. Here, $\boldsymbol{\beta}^{\ast}:=\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$. It is seen that the magnitudes of entries in $\boldsymbol{\beta}^{\ast}$ reflect the feature importance. Motivated by this, we select variables by thresholding a rough estimate of the sparse discriminating direction $\boldsymbol{\beta}^{\ast}$, used solely to guide variable selection without aiming for precise support recovery. Note that if the true clusters $G_{1}^{*}$ and $G_{2}^{*}$ were known, $\boldsymbol{\beta}^{\ast}$ can be accurately estimated by the sample mean difference $\bar{\tilde{\mathbf{X}}}_{G_{1}^{*}}-\bar{\tilde{\mathbf{X}}}_{G_{2}^{*}}$, where $\bar{\tilde{\mathbf{X}}}_{G_{k}^{*}}\in\mathbb{R}^{p}$ stands for the sample mean vector formed by columns of $\tilde{\mathbf{X}}$ in cluster $k=1,2$.
@@ 3.1 Iterative algorithm under known covariance using the innovated transformation
$\hat{S}^{t}:=\biggl{\{}j\in\{1,\ldots,p\}:|\hat{\beta}^{t}_{j}|>\sqrt{2\omega_{jj}n\log(2p)}/\sqrt{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}\biggr{\}}.$ (7)
We provide some intuition on the above proposed variable selection method. Note that if $j\not\in S_{0}$, then $\tilde{X}_{ji}\overset{i.i.d.}{\sim}\mathcal{N}(\tilde{\mu}_{j},\omega_{jj})$, where $\omega_{jj}$ is the $j$th diagonal entry of $\boldsymbol{\Omega}^{\ast}$ and $\tilde{\mu}_{j}$ is the mean value that is invariant to class label. For any data partition $G_{1}\cup G_{2}=[n]$ and $j\not\in S_{0}$, the $j$th entry of the mean difference vector $\bar{\tilde{\mathbf{X}}}_{G_{1}}-\bar{\tilde{\mathbf{X}}}_{G_{2}}$ follows the distribution $\mathcal{N}\bigl{(}0,(w_{jj}n)/(|G_{1}|\cdot|G_{2}|)\bigr{)}$. Then we can derive that for any $\delta>0$,
$\mathbb{P}\biggl{(}\max_{G_{1},G_{2}}\max_{j\in S_{0}^{c}}\left|\sqrt{\frac{|G_{1}|\cdot|G_{2}|}{\omega_{jj}}}(\bar{\tilde{\mathbf{X}}}_{G_{1}}-\bar{\tilde{\mathbf{X}}}_{G_{2}})_{j}\right|\leq\sqrt{2n\log\bigl{(}2(p-|S_{0}|)/\delta\bigr{)}}\biggr{)}\geq 1-\delta,$ (8)
by the Gaussian maximal inequality (see Appendix F.1 for details). We choose a slightly more conservative threshold in (7), following the intuition discussed in the last section that some mild over selection does not hurt clustering.
Clustering step.
At iteration $t$, the clustering step solves the reduced version of SDP-relaxed K-means problem (2), formulated with the selected variables $\hat{S}^{t}$ and using the identity $\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}=\tilde{\mathbf{X}}^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{X}}$. Since this step is reused in unknown covariance scenario, we present it as Algorithm 1. In line 1 of Algorithm 1, $\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}^{\top}\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}$ is different from $(\tilde{\mathbf{X}}^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{X}})_{\hat{S}^{t},\hat{S}^{t}}$, and captures the key sparse clustering signal in $\boldsymbol{\beta}^{\ast}$ (see Appendix E.2 for details).
Algorithm 1 SDP K-means subroutine for $K=2$ : SDPkmeans
@@
2:$\hat{\mathbf{Z}}^{t}\leftarrow\arg\max_{\mathbf{Z}\in\mathbb{R}^{n\times n}}\langle\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}^{\top}\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot},\mathbf{Z}\rangle\newline \hskip 27.0pt\text{s.t.}\quad\mathbf{Z}^{\top}=~{}\mathbf{Z},~{}\mathbf{Z}\succeq~{}0,~{}\mathrm{tr}(\mathbf{Z})=~{}2,\mathbf{Z}\boldsymbol{1}_{n}=~{}\boldsymbol{1}_{n},~{}\mathbf{Z}\geq 0$
3:$\hat{G}^{t}_{1},\hat{G}^{t}_{2}\leftarrow$ spectral clustering result (for example, Han et al.,, 2023) on matrix $\hat{\mathbf{Z}}^{t}$ as an affinity matrix
@@
2:$\hat{G}_{1}^{0},\hat{G}_{2}^{0}\leftarrow$ cluster estimates obtained by running initial clustering (i.e., spectral clustering) on $\mathbf{X}$
3:$\tilde{\mathbf{X}}\leftarrow\boldsymbol{\Sigma}^{-1}\mathbf{X}$triangleright Transform
4:$(\omega_{11},\ldots,\omega_{pp})\leftarrow\operatorname{diag}(\boldsymbol{\Sigma}^{-1})$
@@
6:     $\hat{\boldsymbol{\beta}}^{t}\leftarrow\bigl{(}|\hat{G}^{t-1}_{1}|^{-1}\sum_{i\in\hat{G}^{t-1}_{1}}\tilde{\mathbf{X}}_{i}\bigr{)}-\bigl{(}|\hat{G}^{t-1}_{2}|^{-1}\sum_{i\in\hat{G}^{t-1}_{2}}\tilde{\mathbf{X}}_{i}\bigr{)}$
7:     $\hat{S}^{t}\leftarrow\left\{j\in\{1,\ldots,p\}:|\hat{\beta}^{t}_{j}|>\sqrt{2\omega_{jj}n\log(2p)}/\sqrt{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}\right\}$triangleright Variable selection
8:     $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{SDPkmeans}(\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot},\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}})$triangleright Clustering (Algorithm 1)
@@
3.2 Iterative algorithm under unknown covariance
We now extend Algorithm 2 to the unknown covariance scenario by adapting the estimation idea proposed in Fan and Lv, (2016). The major advantage of our estimation method is that it does not require estimating the full precision matrix $\boldsymbol{\Omega}^{*}$. The main idea is that, at the $t$th iteration, we form an estimate of $\tilde{\mathbf{X}}$ directly based on the current cluster assignments without estimating $\boldsymbol{\Omega}^{*}$, and then calculate the estimate $\hat{S}^{t}$ in (7) using the estimated $\tilde{\mathbf{X}}$.
The direct estimation of $\tilde{\mathbf{X}}$ is done via the Innovated Scalable Efficient Estimation (ISEE) procedure (Fan and Lv,, 2016), originally developed for high-dimensional sparse precision matrix estimation. We modify it to selectively estimate only the required quantities for our algorithm, avoiding full precision matrix recovery, and use its theory to set the threshold for variable selection. Section 3.2.1 summarizes the ISEE procedure adapted as a subroutine in our method, followed by the full algorithm in Section 3.2.2.

### v2

@@ 2 THEORETICAL MOTIVATION
$\text{and}~\mathcal{H}(K):=\bigl\{\mathbf{H}=(H_{i,k})\in\{0,1\}^{n\times K}:\sum_{k=1}^{n}H_{i,k}=1~\text{for all }i=1,\ldots,n\bigr\}.$
For a given clustering algorithm $\hat{\mathbf{H}}:\mathbb{R}^{p\times n}\to\mathcal{H}(K)$, we assess its performance using the worst-case probability of exact cluster recovery over the distribution class $\Theta(s,\Delta,K)$:
$\sup_{\boldsymbol{\theta}\in\Theta(s,\Delta,K)}\min_{\pi\in\mathcal{S}_{K}}\mathbb{P}_{\boldsymbol{\theta}}\bigl(\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H}\bigr),$ (7)
where $\mathcal{S}_{K}$ is the set of label permutations of $[K]$. For brevity, we omit $\min_{\pi\in\mathcal{S}_{K}}$ henceforth.
Given a problem instance $\boldsymbol{\theta}^{\ast}=(\boldsymbol{\mu}_{1}^{\ast},\ldots,\boldsymbol{\mu}_{k}^{\ast},\mathbf{H}^{\ast})\in\Theta(s,\Delta,K)$, the true cluster assignment $\mathbf{H}^{\ast}$ has one-to-one correspondence to the block-diagonal matrix $\mathbf{Z}^{\ast}:=\mathbf{H}^{\ast}\mathbf{B}^{\ast}(\mathbf{H}^{\ast})^{\top}=\sum_{k=1}^{K}|G_{k}^{\ast}|^{-1}\mathbf{1}_{G_{k}^{\ast}}\mathbf{1}_{G_{k}^{\ast}}^{\top},$ up to label permutations. Let the clustering signal (or separation) of a feature subset $S$ be defined as
$\Delta_{S\cap S_{0}}^{2}:=\min_{1\leq k\neq l\leq K}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}.$ (8)
We next define the collection of feature subsets with sufficiently strong signal:
$\mathcal{S}:=\bigl\{S:S\subset[p],|S|\leq\sqrt{p},\Delta_{S\cap S_{0}}^{2}\gtrsim\bar{\Delta}_{S}^{2}\bigr\},$ (9)
$\text{where}~\bar{\Delta}_{S}^{2}:=\sigma^{2}\biggl(\log n+\frac{|S|\log p}{m}+\sqrt{\frac{|S|\log p}{m}}\biggr),$
$\text{and}~m:=2\min_{1\leq k\neq l\leq K}\left\{\biggl(\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr)^{-1}\right\}.$
Theorem 1 gives the separation required for SDP $K$-means restricted to $S$ to achieve exact recovery:
@@
Assume that there exists a universal constant $C_{1}$ such that $m\geq C_{1}n/\log n.$ Also assume $|G_{k}^{\ast}|\geq 2$ for all $k\in[K]$. Let $\hat{\mathbf{Z}}(S)$ be the solution of (5) corresponding to the subset $S\in\mathcal{S}$. Then
$\mathbb{P}\bigl(\hat{\mathbf{Z}}(S)=\mathbf{Z}^{*},\forall S\in\mathcal{S}\bigr)\geq 1-C_{3}K/n,$
where $C_{3}>0$ is some constant.
The proof of Theorem 1 is presented in Appendix C. The stated probability in Theorem 1 holds uniformly over all feature subsets in $\mathcal{S}$. This ensures the validity of iterative algorithms that solve a sequence of SDPs corresponding to the sequence of selected feature subsets: once the algorithm lands on a ‘good’ $S$, that step ensures accurate clustering. The assumption $m\geq C_{1}n/\log n$ prevents overly imbalanced clusters. Such conditions are essential for theoretical guarantees in clustering and are similarly imposed in previous works (Azizyan et al.,, 2013; Cai et al.,, 2019; Chen and Yang,, 2021; Chen and Zhang,, 2024; Jin and Wang,, 2016; Löffler et al.,, 2021; Verzelen and Arias-Castro,, 2017; Giraud and Verzelen,, 2019; Royer,, 2017; Even et al.,, 2025). By definition of $m$ as a harmonic average, the restriction on $m$ implies a restriction on the cluster size $K\lesssim\log n$, which allows a known $K$ to diverge slowly with $n$. This restriction is milder than most of the previous literature, which often requires $K$ bounded above by a constant (Verzelen and Arias-Castro,, 2017; Jin and Wang,, 2016).
Next, we show the optimality of result in Theorem 1.
Theorem 2.
Consider the regime $(s\log p)/n=o(1)$. For $n\geq 2$, no clustering rule can guarantee exact recovery in the worst case over the class
$\overline{\Theta}:=\Theta\bigl(s,C_{4}\sigma^{2}\log n,K\bigr),$
with $C_{4}$ some small positive constant. In other words,
$\inf_{\hat{\mathbf{H}}:\mathbb{R}^{p\times n}\to\mathcal{H}(K)}\sup_{\boldsymbol{\theta}\in\overline{\Theta}}\mathbb{P}_{\boldsymbol{\theta}}\bigl(\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H}\bigr)\geq C_{5},$
with $C_{5}$ a positive constant.
The proof of Theorem 2 is presented in Appendix D. Theorem 2 implies that no clustering rule can provide a worst-case guarantee of exact recovery across all subsets $S$ whose signal strength is of order $\sigma^{2}\log n$. Under the regime of $(|S|\log p)/m\lesssim\log n$, the separation condition in Theorem 1 (i.e., lower bound for $\Delta^{2}_{S\cap S_{0}}$) is of order $\log n$ and thus matches the bound in Theorem 2 up to a constant, demonstrating the minimax optimality of Theorem 1 result.
Theorems 1 and 2 also indicate that, in the sparse regime where $|S_{0}|\ll p$, feature selection can be critical for clustering accuracy. To see this, consider the simple scenario where $K=2$, $|G_{1}^{\ast}|=|G_{2}^{\ast}|$ and $(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})_{S_{0}}=\mu_{0}\mathbf{1}_{|S_{0}|}$. Then Theorems 1 and 2 prove that for any set $S$ satisfying $|S|\lesssim\sqrt{p}$ and $|S|\lesssim(n\log n)/\log p$ , exact recovery is possible if and only if
@@
This condition becomes very stringent if $S$ is a very noisy estimate of $S_{0}$ (i.e., missing too many signal variables). It also suggests that a slight overestimate of $S_{0}$ may not be very damaging. These motivate us to design an algorithm that incorporates some conservative feature selection method into clustering.
3 PROPOSED INNOVATED TRANSFORMATION-BASED ITERATIVE ALGORITHMS
We propose our algorithms under scenarios of known covariance (Section 3.1), which includes the settings of Theorems 1 and 2, and unknown covariance (Section 3.2). The algorithms are presented for $K=2$ for simplicity, although they naturally extend to larger $K$.
3.1 Known Covariance Case
The success of feature selection hinges on signal strength. In the related but different problem of classification, it has been recognized (Fan et al.,, 2013, 2015) that the innovated transformation $\tilde{\mathbf{X}}=\boldsymbol{\Omega}^{\ast}\mathbf{X}$ maximally boosts the signal strength of important features for a general precision matrix that may not be identity matrix. See Fan et al., (2013) for detailed arguments and optimality study for the innovated transformation in Gaussian classification. Our algorithms work with the innovated transformation $\tilde{\mathbf{X}}$ (or its estimate when $\boldsymbol{\Omega}^{\ast}$ unknown, in Section 3.2). It alternates between a feature selection and SDP $K$-Means, detailed below, with the complete procedure summarized in Algorithm 2.
Feature Selection
As argued in Cai et al., (2019), Gaussian clustering is closely related to classification where the optimal rule is the Bayes rule with linear classification boundary $\mathbf{x}^{T}\boldsymbol{\beta}^{\ast}$ for a given test point $\mathbf{x}\in\mathbb{R}^{p}$. Here, $\boldsymbol{\beta}^{\ast}:=\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$. It is seen that the magnitudes of entries in $\boldsymbol{\beta}^{\ast}$ reflect the feature importance. Motivated by this, we select features by thresholding a rough estimate of $\boldsymbol{\beta}^{\ast}$, used solely to guide feature selection without aiming for precise support recovery. Note that if the true clusters $G_{1}^{*}$ and $G_{2}^{*}$ were known, $\boldsymbol{\beta}^{\ast}$ can be accurately estimated by the sample mean difference $\bar{\tilde{\mathbf{X}}}_{G_{1}^{*}}-\bar{\tilde{\mathbf{X}}}_{G_{2}^{*}}$, where $\bar{\tilde{\mathbf{X}}}_{G_{k}^{*}}\in\mathbb{R}^{p}$ stands for the sample mean vector formed by columns of $\tilde{\mathbf{X}}$ in cluster $k=1,2$.
@@ 3.1 Known Covariance Case
$\hat{S}^{t}:=\biggl\{j\in[p]:|\hat{\beta}^{t}_{j}|>\sqrt{\frac{2\omega_{jj}n\log(2p)}{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}}\biggr\}.$ (10)
We provide some intuition on the above proposed feature selection method. Note that if $j\not\in S_{0}$, then $\tilde{X}_{ji}\overset{i.i.d.}{\sim}\mathcal{N}(\tilde{\mu}_{j},\omega_{jj})$, where $\omega_{jj}$ is the $j$th diagonal entry of $\boldsymbol{\Omega}^{\ast}$ and $\tilde{\mu}_{j}$ is the mean value that is invariant to class label. For any data partition $G_{1}\cup G_{2}=[n]$ and $j\not\in S_{0}$, the $j$th entry of the mean difference vector $\bar{\tilde{\mathbf{X}}}_{G_{1}}-\bar{\tilde{\mathbf{X}}}_{G_{2}}$ follows the distribution $\mathcal{N}\bigl(0,(w_{jj}n)/(|G_{1}|\cdot|G_{2}|)\bigr)$. Then we can derive that for any $\delta>0$, with probability at least $1-\delta$,
$\max_{G_{1},G_{2}}\max_{j\in S_{0}^{c}}\biggl|\sqrt{\frac{|G_{1}|\cdot|G_{2}|}{\omega_{jj}}}(\bar{\tilde{\mathbf{X}}}_{G_{1}}-\bar{\tilde{\mathbf{X}}}_{G_{2}})_{j}\biggr|\leq\sqrt{2n\log\bigl(2(p-|S_{0}|)/\delta\bigr)},$ (11)
by the Gaussian maximal inequality (see Appendix F.1 for details). Our threshold in (10) is slightly more conservative than the right-hand side of (11), reflecting the intuition that mild over-selection does not harm clustering, as formalized by Theorems 1 and 2.
Clustering
At iteration $t$, the clustering step solves the reduced version of SDP-relaxed $K$-means problem (4), formulated with the selected variables $\hat{S}^{t}$ and using the identity $\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}=\tilde{\mathbf{X}}^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{X}}$. Since this step is reused in unknown covariance scenario, we present it as Algorithm 1. In line 1 of Algorithm 1, $\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}^{\top}\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}$ is different from $(\tilde{\mathbf{X}}^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{X}})_{\hat{S}^{t},\hat{S}^{t}}$, and captures the key sparse clustering signal in $\boldsymbol{\beta}^{\ast}$. This argument is justified in Appendix E.2 and numerically evaluated in Appendix H.1.
Algorithm 1 SDP $K$-means for $K=2$ : clust
@@
[…]
@@
[…]
@@
6:  $\hat{\boldsymbol{\beta}}^{t}\leftarrow\frac{1}{|\hat{G}^{t-1}_{1}|}\sum_{i\in\hat{G}^{t-1}_{1}}\tilde{\mathbf{X}}_{i}-\frac{1}{|\hat{G}^{t-1}_{2}|}\sum_{i\in\hat{G}^{t-1}_{2}}\tilde{\mathbf{X}}_{i}$
7:  $\hat{S}^{t}\leftarrow\biggl\{j\in[p]:|\hat{\beta}^{t}_{j}|>\sqrt{\frac{2\omega_{jj}n\log(2p)}{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}}\biggr\}$ $\triangleright$ (10)
8:  $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{clust}(\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot},\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}})$ $\triangleright$ Algorithm 1
@@
3.2 Unknown Covariance Case
We now extend Algorithm 2 to the unknown covariance scenario by adapting the estimation idea proposed in Fan and Lv, (2016). The major advantage of our estimation method is that it does not require estimating the full precision matrix $\boldsymbol{\Omega}^{*}$. The main idea is that, at the $t$-th iteration, we form an estimate of $\tilde{\mathbf{X}}$ directly based on the current cluster assignments without estimating $\boldsymbol{\Omega}^{*}$, and then calculate the estimate $\hat{S}^{t}$ in (10) using the estimated $\tilde{\mathbf{X}}$.
The direct estimation of $\tilde{\mathbf{X}}$ is done via the Innovated Scalable Efficient Estimation (ISEE) procedure (Fan and Lv,, 2016), originally developed for high-dimensional sparse precision matrix estimation. We modify it to selectively estimate only the required quantities for our algorithm, avoiding full precision matrix recovery, and use its theory to set the threshold for feature selection. Section 3.2.1 summarizes the ISEE procedure adapted as a subroutine in our method, followed by the full algorithm in Section 3.2.2.